OpenStudy (anonymous):

Help me please Andy has 26 coins made up of quarters and half dollars, and their total value is $11.50. How many quarters does he have?

OpenStudy (lilg132):

x = half dollars y = quarters x + y = 26 x = 26 - y y = 26 - x x(50) + y(25) = 1150 50x + 25y = 1150 50(26 - y) + 25y = 1150 1300 - 50y + 25y = 1150 1300 - 1150 = 50y - 25y 150 = 25y y = 6 50x + 25(6) = 1150 50x + 150 = 1150 50x = 1150 - 150 50x = 1000 x = 20 20 half dollar coins and 6 quarter dollar coins dont forget to click good answer :)

OpenStudy (anonymous):

Why did you use 1150 instead of $11.50?

OpenStudy (lilg132):

it just makes it easier you can use 11.50 if you want and use x(0.50) + y(0.25) instead of x(50) + y(25)

OpenStudy (lilg132):

instead of using decimal points

OpenStudy (anonymous):

oh! thanks :)

OpenStudy (lilg132):

your welcome